Blue rulers cost $1.50 and green rulers cost $4.50. The school purchases total of 60 rulers for the students and spends $129. How many of each type of ruler did she buy?
Answer by ankor@dixie-net.com(22740)

let b = no. of blue rulers
let g = no. of green
:
Blue rulers cost $1.50 and green rulers cost $4.50.
The school purchases total of 60 rulers for the students and spends $129.
How many of each type of ruler did she buy?
:
two equations
b + g = 60
or
b = (60-g)
and
1.50b + 4.50g = 129
replace b with (60-g)
1.5(60-g) + 4.5g = 129
90 - 1.5g + 4.5g = 129
3g = 129 - 90
3g = 39
g = 39/3
g = 13 green rulers
you can find the no. of blue ones
:
Check solutions in the 2nd equation